# ControlDreamer: Stylized 3D Generation with Multi-View ControlNet

### v2

@@ Abstract
Recent advancements in text-to-3D generation have significantly contributed to the automation and democratization of 3D content creation. Building upon these developments, we aim to address the limitations of current methods in generating 3D models with creative geometry and styles. We introduce multi-view ControlNet, a novel depth-aware multi-view diffusion model trained on generated datasets from a carefully curated $100K$ text corpus. Our multi-view ControlNet is then integrated into our two-stage pipeline, ControlDreamer, enabling text-guided generation of stylized 3D models. Additionally, we present a comprehensive benchmark for 3D style editing, encompassing a broad range of subjects, including objects, animals, and characters, to further facilitate diverse 3D generation. Our comparative analysis reveals that this new pipeline outperforms existing text-to-3D methods as evidenced by qualitative comparisons and CLIP score metrics.
Keywords:
text-to-3D, two-stage pipeline, multi-view diffusion model, ControlNet
@@ 1 Introduction
3D content creation is recently gathering attention, particularly in the realms of virtual reality and game development. This evolving landscape has been profoundly impacted by the recent advancement of 2D lifting methods. Dreamfusion’s [29] introduction of score distillation sampling (SDS) has revolutionized and democratized the field, facilitating the generation of 3D models by leveraging large-scale text-to-image diffusion models [35, 33, 1]. This breakthrough has enabled more intuitive and imaginative 3D content creation from user’s textual descriptions. Furthering this evolution, various works show high-resolution text-to-3D generation, employing multiple 3D representations [21, 4, 5] or variants of score distillation sampling [42, 11].
However, current 2D lifting methods face the Janus Problem of generating multiple faces from different angles due to a lack of 3D knowledge. MVDream [39] addresses this by adapting Stable Diffusion’s [33] 2D self-attention to 3D and jointly training with multi-view images from the Objaverse [8] and LAION datasets [36]. For instance, prompts like ‘Captain America’ tend to yield 3D model of a shield rather than the character, likely due to dataset biases and the mode-seeking nature of score distillation sampling [29]. This leads to the generation of mode geometries learned from the dataset, which limits the creation of customized 3D models, such as a bulky Captain America (Fig. 1).
@@
Figure 1: Comparing text-to-3D pipeline of our ControlDreamer and MVDream. The right side showcases MVDream’s output, which demonstrates a vulnerability to geometry biases learned during pre-training, often resulting in (a) incomplete models like only generating Captain America’s shield or a lioness’s head or (b) stereotypical geometries related to the prompts. In contrast, the left side displays how ControlDreamer, through its two-stage pipeline, effectively overcomes geometry biases, enabling unique combinations of geometry and style in 3D models, even for counterfactual generation.
@@
These challenges motivated us to study a two-stage approach given two text prompts. Initially, we generate a neural radiance field (NeRF) [23, 24] from a geometry prompt to sculpt geometric structures, followed by converting it to a mesh using DMTet algorithm [38] and refining both geometry and texture using a style prompt. In the second stage, we found that pre-trained MVDream model encounters difficulties in aligning textures with the initially generated geometrical structures, resulting in noticeable artifacts as represented in Fig. 2 (b). To address this, we present MV-ControlNet, a depth-aware multi-view diffusion model. We first generate a paired dataset of text and multi-view images using GPT-4 filtered texts and MVDream, followed by training a depth conditioner for the frozen MVDream, according to the principle of ControlNet [43]. Obtaining MV-ControlNet is cost-effective, requiring only 1.5 days on dataset generation and 1.5 days on training, on four A40 GPUs (Note, MVDream required 3 days on 32 A100 GPUs for model training).
@@
In evaluations conducted on our new benchmark, which consists of pairs of geometry and style prompts, ControlDreamer demonstrates superior performance in generating styles on 3D models compared to existing two-stage pipelines. By using CLIP scores and human assessments, we find that our depth-aware MV-ControlNet outperforms normals and edges-aware variants in rendering detailed textures and geometries. These findings confirm ControlDreamer’s ability to produce high-quality 3D models that closely match the provided textual descriptions. For more detailed results and comparisons, please visit our project page.
Our contributions can be summarized as follows:
  • We introduce ControlDreamer, a two-stage pipeline for 3D model generation, combining NeRF creation from geometry prompts and mesh refinement with style prompts.
  • The training of MV-ControlNet is more practical and resource-efficient, offering improvements in time and resource usage compared to prior models.
  • Our MV-ControlNet adeptly aligns diverse geometries and styles, elevating the generation of view-consistent geometry and styles from various texts.
@@ 2.1 Adding Conditions to Text-to-Image Models
Diffusion models like Stable Diffusion [33], Imagen [35], and eDiff-I [1], trained on billions of image-text pairs, have shown remarkable capabilities in generating high-quality images from text. ControlNet [43] introduces a method for adding conditions to pre-trained text-to-image models by using a trainable copy alongside the original frozen model for specific tasks like depth-to-image. Instruct-Pix2Pix [2] uses GPT-3 [3] and Stable Diffusion to create a training dataset for an image editing model that follows diverse editing instructions.
Figure 2: Results from generating various stylized 3D models using style prompts, based on a source NeRF created from a geometry prompt. Our ControlDreamer (a) aligns various styles on the source geometry, while the baseline (b) suffers artifacts due to misalignment of style and geometry.
@@ 2.2 Text-to-3D Generation
Training a generative 3D model using large-scale 3D datasets is highly challenging due to the difficulty and expense of collecting 3D data compared to images. DreamFusion [29] introduces the score distillation sampling (SDS), enabling the optimization of an implicit 3D model [23, 24] without any 3D dataset by utilizing a pre-trained 2D diffusion model as a prior. Despite these advantages, DreamFusion has the drawback of generating scenes very slowly and producing lower-quality results. To overcome these limitations, several methods are proposed involving primary optimization for NeRF and secondary optimization for the extracted mesh [21, 5], along with improvements for SDS [42, 15].
Table 1: Previous text-to-3D generation scennarios. Our method is the first two-stage pipeline designed for multi-view consistent 3D generation. Notably, it offers convenience by eliminating the need for annotations like bounding boxes and masks.
Method 3D Rep. Multi-Stage Annot.-Free Multi-View
Vox-E [37] Voxel
PG3D [6] NeRF
TEXTure [32] Texture
Text2tex [4] Texture
ProlificDreamer [42] NeRF
Fantasia3D [5] DMTet
Magic3D [21] NeRF&\&&DMTet
MVDream [39] NeRF
Ours NeRF&\&&DMTet
2.3 Multi-view Diffusion Nodel
In contrast to prior 2D lifting-based methodologies, recent advancements have introduced multi-view aware training for text-to-image diffusion models. MV-diffusion [40] facilitates the generation of multi-view panoramic images and depth-to-image conversions using given text prompts. This approach introduces a novel attention module to account for pixel-to-pixel correspondences, ensuring coherence across different views. MVDream [39] integrates a 3D self-attention module and fine-tunes the diffusion model through joint training on large-scale 2D and 3D datasets. As a result, MVDream avoids the Janus Problem and Content Drifting Problem, consistently producing 3D models that align accurately with the provided texts.
@@ 2.4 Text-Guided 3D Editing
In text-to-3D generation, various innovative methods have been proposed. Vox-E [37] edits voxel grids with SDS, maintaining geometry and appearance. Instruct-NeRF2NeRF [11] iteratively updates source NeRFs using Instruct-Pix2Pix [2]. Text2tex [4] assumes a perfect mesh, applying textures without mesh refinement flexibility. Magic3D [21] offers a coarse-to-fine pipeline for high-resolution 3D models, starting with NeRF and converting to deformable meshes. Fantasia3D [5] divides its process into geometry and appearance stages, beginning with a 3D ellipsoid and refining with DMTet. Progressive3D [6] decomposes complex prompts for localized editing, focusing on user-specified regions.
In Table 1, we outline a key gap in existing research: a lack of methods combining multi-stage, annotation-free, and multi-view capabilities for 3D generation. We found that current models, especially in DMTet texture refinement, work best when geometry and style prompts match. Furthermore, methods using 2D models struggle with multi-view consistency. Addressing this, we integrated multi-view features from MVDream in our ControlDreamer, enabling high-quality, differentiated prompt-based 3D generation.
@@ 3 Method
Our ControlDreamer pipeline involves a two-stage process: first, training a coarse-grained NeRF with a geometry prompt and converting it into a textured mesh using DMTet, and then creating a fine-grained stylized 3D model with a style prompt. We illustrate our pipeline in Fig. 3 and provide detailed explanations in Sec. 3.1. Then, we emphasize a depth-aware, multi-view approach in the second stage and elaborate on the Multi-View (MV) ControlNet’s training scheme in Sec. 3.2.
Figure 3: Illustration of ControlDreamer. (Left) Using a geometry prompt and MVDream, we create a multi-view consistent NeRF, with MVDream’s 3D self-attention ensuring consistency. (Right) The NeRF is transformed into a mesh via DMTet and undergoes style generation using our MV-ControlNet, which combines a trainable copy (red) and zero convolution (white). MV-ControlNet is designed to understand geometry from the multi-view depths, enabling effective style generation.
@@ 3.1.1 Geometry Stage
Here, we generate a NeRF from a given geometry prompt. The style of this NeRF will subsequently be modified in the second stage. We optimize the NeRF by distilling probability density [26] from the pre-trained MVDream [39]. This optimization is achieved through score distillation sampling (SDS) [29], where the diffusion model directly computes the loss gradient as shown in the following equation:
$\nabla_{\theta}\mathcal{L}_{\text{SDS}}\triangleq\mathbb{E}_{t,\epsilon}\left[w(t)\left(\hat{\epsilon}_{\phi}(z_{t};y_{geo},c,t)-\epsilon\right)\frac{\partial z}{\partial x}\frac{\partial x}{\partial\theta}\right],$ (1)
where $t\sim U(t_{\text{min}},t_{\text{max}})$, $w(t)$ denotes weighting function, $\hat{\epsilon}_{\phi}(z_{t};y,t)$ represents a guided epsilon prediction [14], $y_{geo}$ is the geometry text prompt, $c$ represents camera parameter, $z$ refers to VAE latents [18, 33] of rendered multi-view images $x$, and $\theta$ is the parameter of 3D model. Note that there is no backpropagation through the diffusion model.
To leverage the multi-view consistency of MVDream, we render four views from randomly selected camera parameters $c$ in each iteration of optimization. These parameters are uniformly sampled from a range of view angles, all maintaining a consistent elevation to ensure a diverse yet coherent range of perspectives. In further detail, we incorporate lighting strategies such as point lighting [29], orientation loss [29], and soft shading [21]. Additionally, we enhance the background with random color augmentation and employ a fixed negative prompt to avoid low-quality 3D models. The timestep range $[t_{\text{min}},t_{\text{max}}]$ is linearly annealed during the optimization process. To prevent color saturation, we apply the $x_{0}$-reconstruction loss, which is the epsilon reconstruction loss weighted by the signal-to-noise ratio (SNR) [17], thus $w(t)=\text{SNR}(t)$. The text prompt $y_{geo}$ is embedded using the CLIP ViT-L text encoder [30].

### v3

@@ Abstract
Recent advancements in text-to-3D generation have significantly contributed to the automation and democratization of 3D content creation. Building upon these developments, we aim to address the limitations of current methods in blending geometries and styles in text-to-3D generation. We introduce multi-view ControlNet, a novel depth-aware multi-view diffusion model trained on generated datasets from a carefully curated text corpus. Our multi-view ControlNet is then integrated into our two-stage pipeline, ControlDreamer, enabling text-guided generation of stylized 3D models. Additionally, we present a comprehensive benchmark for 3D style editing, encompassing a broad range of subjects, including objects, animals, and characters, to further facilitate research on diverse 3D generation. Our comparative analysis reveals that this new pipeline outperforms existing text-to-3D methods as evidenced by human evaluations and CLIP score metrics. Project page: https://controldreamer.github.io
@@ 1 Introduction
3D content creation has recently been gathering attention, particularly in the realms of virtual reality and game development. This evolving landscape has been profoundly impacted by the recent advancement of 2D lifting methods. Dreamfusion’s [Poole et al.(2023)Poole, Jain, Barron, and Mildenhall] introduction of score distillation sampling (SDS) has revolutionized and democratized the field, facilitating the generation of 3D models by leveraging large-scale text-to-image diffusion models [Saharia et al.(2022)Saharia, Chan, Saxena, Li, Whang, Denton, Ghasemipour, Gontijo Lopes, Karagol Ayan, Salimans, et al., Rombach et al.(2022)Rombach, Blattmann, Lorenz, Esser, and Ommer, Balaji et al.(2022)Balaji, Nah, Huang, Vahdat, Song, Kreis, Aittala, Aila, Laine, Catanzaro, et al.]. This breakthrough has enabled more intuitive and imaginative 3D content creation from users’ textual descriptions. Furthering this evolution, various works show high-resolution text-to-3D generation, employing multiple 3D representations [Lin et al.(2023)Lin, Gao, Tang, Takikawa, Zeng, Huang, Kreis, Fidler, Liu, and Lin, Chen et al.(2023a)Chen, Siddiqui, Lee, Tulyakov, and Nießner, Chen et al.(2023b)Chen, Chen, Jiao, and Jia] or variants of score distillation sampling [Wang et al.(2023)Wang, Lu, Wang, Bao, Li, Su, and Zhu, Haque et al.(2023)Haque, Tancik, Efros, Holynski, and Kanazawa]. One such approach is MVDream [Shi et al.(2023)Shi, Wang, Ye, Long, Li, and Yang], a model that extends the capabilities of Stable Diffusion [Rombach et al.(2022)Rombach, Blattmann, Lorenz, Esser, and Ommer], a well-known text-to-image diffusion model, by adapting it on the Objaverse dataset [Deitke et al.(2023)Deitke, Schwenk, Salvador, Weihs, Michel, VanderBilt, Schmidt, Ehsani, Kembhavi, and Farhadi], which consists of multi-view images of 3D objects.
However, we observe that when attempting to generate imaginative 3D models using MVDream, it fails to seamlessly blend attributes such as geometry and style. For instance, when given a prompt like ‘Captain America,’ the model tends to produce a 3D model of a shield rather than the character itself. This issue likely arises due to dataset biases and the mode-seeking nature of score distillation sampling [Poole et al.(2023)Poole, Jain, Barron, and Mildenhall], which causes the model to generate geometries that reflect dominant modes in the dataset, rather than customized or more nuanced 3D models, such as a bulky version of Captain America (Fig. 1).
@@
Figure 1: Comparing text-to-3D pipeline of our ControlDreamer and MVDream. On the right, MVDream’s output reveals a vulnerability to pre-training geometry biases, often producing (a) unintended results such as shields or (b) stereotypical geometries related to prompts. On the left, ControlDreamer overcomes these biases, enabling unique combinations of geometry and style, even facilitating counterfactual generation in 3D models.
@@
These challenges motivated us to decompose the text-to-3D pipeline into a two-stage approach given two text prompts. Initially, we generate a neural radiance field (NeRF) [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng, Müller et al.(2022)Müller, Evans, Schied, and Keller] from a geometry prompt to sculpt geometric structures, followed by converting it to a mesh using DMTet algorithm [Shen et al.(2021)Shen, Gao, Yin, Liu, and Fidler] and refining both geometry and texture using a style prompt. To address this, we present MV-ControlNet, a depth-aware multi-view diffusion model. We first generate a paired dataset of text and multi-view images using GPT-4 [OpenAI(2023)] filtered texts and MVDream, followed by training a depth conditioner for the frozen MVDream, according to the principle of ControlNet [Zhang et al.(2023)Zhang, Rao, and Agrawala].
@@
In evaluations on our new benchmark, which consists of pairs of geometry and style prompts, ControlDreamer shows superior performance in generating styles on 3D models compared to existing two-stage pipelines. By using CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al.] scores and human assessments, we find that our depth-aware MV-ControlNet outperforms normal and edge-aware variants in rendering detailed 3D models. These findings confirm ControlDreamer’s ability to produce high-quality 3D models that closely match the provided textual descriptions.
Table 1: Comparison to previous text-to-3D generation methods. Our method is the first two-stage pipeline designed for multi-view consistent 3D generation. Notably, it offers convenience by eliminating the need for annotations like bounding boxes and masks.
Method 3D Rep. Multi-Stage Annot.-Free Multi-View
ProlificDreamer [Wang et al.(2023)Wang, Lu, Wang, Bao, Li, Su, and Zhu] NeRF
Fantasia3D [Chen et al.(2023b)Chen, Chen, Jiao, and Jia] DMTet
Magic3D [Lin et al.(2023)Lin, Gao, Tang, Takikawa, Zeng, Huang, Kreis, Fidler, Liu, and Lin] NeRF&\&&DMTet
MVDream [Shi et al.(2023)Shi, Wang, Ye, Long, Li, and Yang] NeRF
Ours NeRF&\&&DMTet
@@ 2.1 Adding Conditions to Text-to-Image Models
Diffusion models like Stable Diffusion [Rombach et al.(2022)Rombach, Blattmann, Lorenz, Esser, and Ommer], trained on billions of image-text pairs [Schuhmann et al.(2022)Schuhmann, Beaumont, Vencu, Gordon, Wightman, Cherti, Coombes, Katta, Mullis, Wortsman, et al.], have shown remarkable capabilities in generating high-quality images from text or in removing harmful degradations from images [Oh et al.(2024)Oh, Lee, Choi, Jung, Hwang, and Yoon]. ControlNet [Zhang et al.(2023)Zhang, Rao, and Agrawala] introduces a method for adding conditions to Stable Diffusion by using a trainable copy alongside the original frozen model to perform specific tasks, such as depth-to-image conversion.
@@ 2.2 Text-to-3D Generation
While public datasets of 2D images are abundant, dataset of 3D assets are relatively scarce ($\sim$80K in Objaverse [Deitke et al.(2023)Deitke, Schwenk, Salvador, Weihs, Michel, VanderBilt, Schmidt, Ehsani, Kembhavi, and Farhadi]). Consequently, leveraging pre-trained 2D diffusion models to create diverse, user-customized 3D models continues to be a demanding task. To overcome such difficulties, DreamFusion [Poole et al.(2023)Poole, Jain, Barron, and Mildenhall] introduces score distillation sampling (SDS) to enable the optimization of an implicit 3D model [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng, Müller et al.(2022)Müller, Evans, Schied, and Keller] without any 3D dataset, utilizing a pre-trained 2D diffusion model as a prior. Despite these advantages, DreamFusion suffers from slow scene generation and lower-quality results. To overcome these limitations, several methods are proposed involving primary optimization for NeRF and secondary optimization for the extracted mesh [Lin et al.(2023)Lin, Gao, Tang, Takikawa, Zeng, Huang, Kreis, Fidler, Liu, and Lin, Chen et al.(2023b)Chen, Chen, Jiao, and Jia], along with improvements for SDS [Wang et al.(2023)Wang, Lu, Wang, Bao, Li, Su, and Zhu, Katzir et al.(2023)Katzir, Patashnik, Cohen-Or, and Lischinski, Metzer et al.(2023)Metzer, Richardson, Patashnik, Giryes, and Cohen-Or, Chen et al.(2024a)Chen, Yang, Yang, Feng, Fu, Foo, Lin, and Liu]. In contrast to prior 2D lifting-based methodologies, MVDream [Shi et al.(2023)Shi, Wang, Ye, Long, Li, and Yang] fine-tunes a text-to-image diffusion model [Rombach et al.(2022)Rombach, Blattmann, Lorenz, Esser, and Ommer] through joint training on large-scale 2D and 3D datasets. Consequently, MVDream circumvents the Janus Problem and the Content Drifting Problem, consistently generating 3D models that align accurately with the given texts. Concurrent works have developed large-scale 3D generative models [Voleti et al.(2024)Voleti, Yao, Boss, Letts, Pankratz, Tochilkin, Laforte, Rombach, and Jampani, Boss et al.(2024)Boss, Huang, Vasishta, and Jampani]. However, these models primarily rely on image inputs, making them unsuitable for our objective of controlling both the geometry and style of 3D models through text descriptions.
Figure 2: Main results. Our generation process begins by generating a coarse-grained geometry, followed by creating a fine-grained stylized 3D model using a style prompt.
@@ 2.3 Text-Guided 3D Editing
In text-to-3D generation, various innovative methods have been proposed. Magic3D [Lin et al.(2023)Lin, Gao, Tang, Takikawa, Zeng, Huang, Kreis, Fidler, Liu, and Lin] offers a coarse-to-fine pipeline for high-resolution 3D models, starting with NeRF and converting to deformable meshes. Fantasia3D [Chen et al.(2023b)Chen, Chen, Jiao, and Jia] divides its process into geometry and appearance stages, beginning with a 3D ellipsoid and refining with DMTet. TEXTure [Richardson et al.(2023)Richardson, Metzer, Alaluf, Giryes, and Cohen-Or] handles a similar task of generating textures on top of geometry, but unlike our approach, it does not allow for changes in the geometry. IT3D [Chen et al.(2024b)Chen, Zhang, Yang, Cai, Yu, Yang, and Lin] refines initial rendered views using a 2D ControlNet [Zhang et al.(2023)Zhang, Rao, and Agrawala] and ensures multi-view consistency through a GAN loss, whereas we use a single multi-view diffusion model. In Table 1, we outline a key gap in existing research: a lack of methods combining multi-stage, annotation-free, and multi-view capabilities for 3D generation. We found that current methods, especially in DMTet texture refinement, work well only when geometry and style prompts match. Furthermore, methods using 2D models struggle with multi-view consistency. Our ControlDreamer generates multi-view consistent 3D models by combining geometry and style specified by text prompts.
Figure 3: Illustration of ControlDreamer. (Left) Starting with a geometry prompt, we use MVDream to generate a NeRF, ensuring consistency through 3D self-attention. (Right) The NeRF is converted into a mesh via DMTet, followed by style generation through our MV-ControlNet, which integrates a trainable copy (red) and employs zero-initialized convolutions (white). MV-ControlNet is designed to understand geometry using multi-view depth.
@@ 3 Method
Our ControlDreamer pipeline involves a two-stage process: first, training a coarse-grained NeRF [Mildenhall et al.(2021)Mildenhall, Srinivasan, Tancik, Barron, Ramamoorthi, and Ng, Müller et al.(2022)Müller, Evans, Schied, and Keller] with a geometry prompt and then creating a fine-grained stylized textured mesh [Shen et al.(2021)Shen, Gao, Yin, Liu, and Fidler] with a style prompt. We illustrate our pipeline in Fig. 3 and provide details in Sec. 3.1. Then, we emphasize a depth-aware, multi-view approach in the second stage and elaborate on the Multi-View (MV) ControlNet’s training scheme in Sec. 3.2.
@@ 3.1.1 Geometry Stage
Here, we generate a NeRF from a given geometry prompt. The style of this NeRF is subsequently modified in the second stage. We optimize the NeRF by distilling probability density [Oord et al.(2018)Oord, Li, Babuschkin, Simonyan, Vinyals, Kavukcuoglu, Driessche, Lockhart, Cobo, Stimberg, et al.] from the pre-trained MVDream [Shi et al.(2023)Shi, Wang, Ye, Long, Li, and Yang]. This optimization is achieved through score distillation sampling (SDS) [Poole et al.(2023)Poole, Jain, Barron, and Mildenhall], where the diffusion model directly computes the loss gradient as shown in the following equation:
$\nabla_{\theta}\mathcal{L}_{\text{SDS}}\triangleq\mathbb{E}_{t,\epsilon}\left[w(t)\left(\hat{\epsilon}_{\phi}(z_{t};y_{\text{geo}},c,t)-\epsilon\right)\frac{\partial z}{\partial x}\frac{\partial x}{\partial\theta}\right],$ (1)
where $t\sim U(t_{\text{min}},t_{\text{max}})$, $w(t)$ denotes the weighting function, $\hat{\epsilon}_{\phi}(z_{t};y,t)$ represents a guided epsilon prediction [Ho and Salimans(2021)], $y_{\text{geo}}$ is the geometry text prompt, $c$ represents the camera parameters, $z$ refers to the VAE latents [Kingma and Welling(2013), Rombach et al.(2022)Rombach, Blattmann, Lorenz, Esser, and Ommer] of rendered multi-view images $x$, and $\theta$ is the parameter of 3D model. Note that there is no backpropagation through the diffusion model.
To leverage the multi-view consistency of MVDream, we render four views using randomly selected camera parameters $c$ in each iteration of optimization. These parameters are uniformly sampled from a range of view angles, each maintaining consistent elevation. In further detail, we incorporate point lighting [Poole et al.(2023)Poole, Jain, Barron, and Mildenhall], orientation loss [Poole et al.(2023)Poole, Jain, Barron, and Mildenhall], and soft shading [Lin et al.(2023)Lin, Gao, Tang, Takikawa, Zeng, Huang, Kreis, Fidler, Liu, and Lin]. The timestep range $[t_{\text{min}},t_{\text{max}}]$ is linearly annealed during the optimization process. To prevent color saturation, we apply the $x_{0}$-reconstruction loss, which is the epsilon reconstruction loss weighted by the signal-to-noise ratio (SNR) [Kingma et al.(2021)Kingma, Salimans, Poole, and Ho], thus $w(t)=\text{SNR}(t)$.